Determine the value of the arbitrary constant of the antriderivative of given the initial value

x = 7.03 and y = 8.75 . (Use 2 decimal places)


1
Expert's answer
2021-03-12T01:06:21-0500

Antriderivative of 1 is x+C.

So 8.75=7.03+C,

C=8.75-7.03=1.72